OpenStudy (anonymous):

What is the value of the expression 7^–3? A. –21 B. -1/343 C. 1/343 D. 343

OpenStudy (anonymous):

B

OpenStudy (anonymous):

Thank you so much if you have time could you help with another?

OpenStudy (anonymous):

alrighty mr. :P

OpenStudy (anonymous):

yolo

OpenStudy (anonymous):

geo get out >:O

OpenStudy (anonymous):

no never lelelelelelelele

OpenStudy (anonymous):

What is the value of the expression 1/3^5 A. 243 B. 15 C 1/15 D 1/243

OpenStudy (anonymous):

D

OpenStudy (anonymous):

yaeh i belive that is right

OpenStudy (anonymous):

Thanks a million i have i think 3 left What is the value of the expression 3/3^2 A. 27 B. 18 C. 1/2 D. 1/3

OpenStudy (anonymous):

D

OpenStudy (anonymous):

What value of k solves the equation? K^4=1/256 A. –64 B. 8 C. 4 D. 64

OpenStudy (anonymous):

d

OpenStudy (anonymous):

2^n=1/8 What value of n solves the equation? A. –3 B. –4 C. 3 D. 4

OpenStudy (anonymous):

@Official_Hotboy

OpenStudy (anonymous):

it is b

OpenStudy (anonymous):

thanks~!

OpenStudy (anonymous):

no its D